# Physics-Informed Singular-Value Learning for Cross-Covariances Forecasting in Financial Markets

### v1

@@ Abstract
A new wave of work on covariance cleaning and nonlinear shrinkage has delivered asymptotically optimal analytical solutions for large covariance matrices. Building on this progress, these ideas have been generalized to empirical cross–covariance matrices, whose singular-value shrinkage characterizes comovements between one set of assets and another. Existing analytical cross-covariance cleaners are derived under strong stationarity and large-sample assumptions, and they typically rely on mesoscopic regularity conditions such as bounded spectra; macroscopic common modes (e.g., a global market factor) violate these conditions. When applied to real equity returns, where dependence structures drift over time and global modes are prominent, we find that these theoretically optimal formulas do not translate into robust out-of-sample performance. We address this gap by designing a random-matrix-inspired neural architecture that operates in the empirical singular-vector basis and learns a nonlinear mapping from empirical singular values to their corresponding cleaned values. By construction, the network can recover the analytical solution as a special case, yet it remains flexible enough to adapt to non-stationary dynamics and mode-driven distortions. Trained on a long history of equity returns, the proposed method achieves a more favorable bias–variance trade-off than purely analytical cleaners and delivers systematically lower out-of-sample cross–covariance prediction errors. Our results demonstrate that combining random-matrix theory with machine learning makes asymptotic theories practically effective in realistic time-varying markets.
@@ 1 Introduction
Related denoising results exist for additive-noise models [17, 11], but sample (cross-)covariances arise from products of observations and therefore call for a multiplicative-noise cleaning framework. To the best of our knowledge, the only available analytical characterization of an asymptotically optimal singular-value shrinkage for empirical cross-covariance matrices under multiplicative noise is due to Benaych-Georges, Bouchaud, and Potters (BBP) [2]. In the high-dimensional asymptotic regime, they derive a Rotationally Invariant Estimator (RIE) that preserves the empirical singular vectors of $\widehat{\mathbf{\Sigma}}_{XY}$ while shrinking its singular values so as to minimize the asymptotic Mean Squared Error (MSE) relative to the true cross-covariance $\mathbf{\Sigma}_{XY}$. Their result is presented in two versions: a baseline setting with identity marginals ($\mathbf{\Sigma}_{XX}=\mathbf{I}_{n_{x}}$ and $\mathbf{\Sigma}_{YY}=\mathbf{I}_{n_{y}}$), which also covers the practically relevant case of whitened variables, and a more general setting that incorporates non-trivial marginal covariances into the optimal correction. The theory is validated through simulations and controlled benchmarks, but it does not consider real financial data and relies on strong stationarity assumptions that are rarely met in practice.
@@
In this work, we introduce a physics-informed learning framework for cross–covariance estimation that preserves the rotational invariances of RMT while operating beyond the stationarity and boundedness regimes required by existing analytical shrinkage. The key idea is to parameterize the cleaned estimator directly in the empirical singular-vector basis and to learn a dimension-agnostic nonlinear map from empirical singular values and marginal projections to cleaned singular values, so that the BBP correction is recovered as a limiting case, but additional degrees of freedom can accommodate macroscopic modes and time variation. Empirically, we observe that the analytically optimal BBP shrinkage can lose robustness on equity returns as the universe size increases in the presence of a dominant market mode, whereas the proposed symmetry-preserving estimator remains stable and yields lower OOS reconstruction error, including in OOS matrix sizes.
@@ 3 Validation on Synthetic and Market Data
Panel A: Finite-Rank Spiked Models (varying spike % $\xi$)
$\xi=40\%$ $2.00\times 10^{-3}$ $2.46\times 10^{-4}$ $\mathbf{2.46\times 10^{-4}}$ $2.53\times 10^{-4}$
$\xi=30\%$ $2.00\times 10^{-3}$ $1.92\times 10^{-4}$ $\mathbf{1.92\times 10^{-4}}$ $1.95\times 10^{-4}$
$\xi=20\%$ $2.00\times 10^{-3}$ $\mathbf{1.34\times 10^{-4}}$ $\mathbf{1.33\times 10^{-4}}$ $1.34\times 10^{-4}$
$\xi=10\%$ $2.00\times 10^{-3}$ $\mathbf{7.03\times 10^{-5}}$ $\mathbf{7.00\times 10^{-5}}$ $7.12\times 10^{-5}$
$\xi=0\%$ $2.00\times 10^{-3}$ $\mathbf{3.73\times 10^{-6}}$ $3.92\times 10^{-6}$ $9.02\times 10^{-6}$
@@
Panel B: Heavy-Tailed Bulk Models (varying tail exponent $\alpha$)
[…]
$\alpha=1.5$ $2.00\times 10^{-3}$ $\mathbf{6.19\times 10^{-4}}$ $6.20\times 10^{-4}$ $\mathbf{6.18\times 10^{-4}}$
@@
Panel C: White Heavy-Tailed Models (varying tail exponent $\alpha$)
Gaussian $2.00\times 10^{-3}$ $\mathbf{5.63\times 10^{-4}}$ $5.64\times 10^{-4}$ $5.64\times 10^{-4}$
$\alpha=5.0$ $2.00\times 10^{-3}$ $\mathbf{5.63\times 10^{-4}}$ $5.65\times 10^{-4}$ $5.64\times 10^{-4}$
$\alpha=2.5$ $2.00\times 10^{-3}$ $\mathbf{2.50\times 10^{-4}}$ $\mathbf{2.50\times 10^{-4}}$ $\mathbf{2.51\times 10^{-4}}$
$\alpha=1.5$ $2.00\times 10^{-3}$ $\mathbf{3.92\times 10^{-5}}$ $\mathbf{3.90\times 10^{-5}}$ $4.14\times 10^{-5}$
@@
Panel D: Gaussian Bulk Models with Common Mode (varying mode $m$)
$m=0.5$ $1.09\times 10^{-3}$ $7.90\times 10^{-2}$ $\mathbf{7.45\times 10^{-4}}$ $\mathbf{7.42\times 10^{-4}}$
$m=0.4$ $1.42\times 10^{-3}$ $3.29\times 10^{-2}$ $\mathbf{9.25\times 10^{-4}}$ $\mathbf{9.18\times 10^{-4}}$
$m=0.3$ $1.66\times 10^{-3}$ $9.89\times 10^{-3}$ $\mathbf{9.83\times 10^{-4}}$ $\mathbf{9.84\times 10^{-4}}$
$m=0.2$ $1.85\times 10^{-3}$ $2.23\times 10^{-3}$ $\mathbf{9.84\times 10^{-4}}$ $\mathbf{9.79\times 10^{-4}}$
$m=0.1$ $1.96\times 10^{-3}$ $8.85\times 10^{-4}$ $\mathbf{8.52\times 10^{-4}}$ $\mathbf{8.56\times 10^{-4}}$
@@
Table 1: Summary of the average MSE for different estimators across three synthetic regimes over 1000 simulations. Bold marks the lowest MSE; ties are bold when a 95% percentile $100,000$ copies bootstrap cannot distinguish them. Panel A varies the spike percentage $\xi$ (Finite-Rank). Panel B and C vary the tail exponent $\alpha$ (Heavy-Tailed Bulk and White models), where a lower $\alpha$ implies heavier tails. Panel D is a variation of Panel B (Gaussian) with a one-factor mode injected. All simulations are performed with $n_{x}=200$, $n_{y}=350$, $\Delta t_{\textrm{in}}=500$ [2].
@@
Figure 2: OOS MSE for cross-correlation estimators on financial data covering 2017–2024. For each test window, the estimator is trained on an expanding sample from 1995 up to the year immediately preceding the test period, and the reported value averages 1,000 independent runs for each out-of-sample year. The figures report also the Oracle (red dashed line), which is the lower achievable MSE by a RIE (see eq. (35)). The left panels preserve the chronological order between in-sample and out-of-sample windows, whereas the right panels shuffle dates. The upper panels varies the total number of assests for $\nu=0.25$, the lower panels varies the relative dimension $\nu$ for $n=1000$. The error bars indicate a $95\%$ percentile $100,000$ copies bootstrap confidence interval.
@@
Figure 2 (upper left) reports the OOS MSE as a function of the total number of assets $n$ for a fixed in-sample window ($\Delta t_{\textrm{in}}=500$). The Maximum-Likelihood Estimator (MLE, yellow) provides the unfiltered baseline. Across both the $n$- and $\nu$-sweeps, the proposed NN estimator (blue) achieves the lowest error. For small universes (e.g., $n=250$), its performance is not statistically distinguishable from BBP (gray). However, as $n$ increases, BBP deteriorates sharply and can even underperform MLE, indicating that the analytical shrinkage can become harmful in this regime. We attribute this failure to the presence of a macroscopic market mode in real returns, which violates the boundedness conditions underlying BBP. This interpretation is supported by two additional diagnostics. First, the numerical CV estimator is not affected by the same degradation and remains comparable to the NN. Second, in Appendix 8.3 we show that pre-processing the data by removing the mode, operationally defined as the daily cross-sectional mean return, eliminates the gross of the BBP underperformance. Finally, when we shuffle the time series to break potential non-stationarity between in-sample and out-of-sample periods, BBP remains unstable whereas CV continues to track the NN, further indicating that the dominant driver is the mode rather than temporal non-stationarity, see Fig. 2 (right) . Notably, the NN performance remains robust even for $n>500$, which constitutes a fully out-of-distribution regime for the model. This region lies beyond the matrix sizes observed during training ($n\leq 500$), and it also corresponds to a different historical period than the data used to fit the network. To further probe this extrapolative regime, in the lower panel of Fig. 2 (lower) we unpack the $n=1000$ case by varying the relative dimension $\nu$, and we find that the NN gain remains stable across the full range of $\nu$, indicating that its advantage is not concentrated in a specific configuration.
@@
Our working hypothesis is that this behavior is precisely the one targeted by the proposed architecture. In stationary settings, the network should primarily act as a noise-cleaning operator on the empirical spectrum. In contrast, when applied to non-stationary financial data, it can interpolate between denoising and a learned predictive correction that exploits persistent structure beyond sampling noise. To confirm that this gain is derived from capturing genuine temporal signals (non-stationarity), we replicate the analysis on temporally shuffled data. Shuffling destroys the serial correlations responsible for the predictive edge, creating a purely stationary environment. Figures 2 (right) illustrate that performance degradation of BBP persists even in the shuffled regime, confirming the underlying structural driver. As expected, the CV estimator does not degrade and instead matches the NN performance indicating that in a stationary regime the NN learn how to denoise only sample noise. This interpretation is further supported by the feasibility diagnostic reported in Appendix 8.4.

### v2

@@ Abstract
A new wave of work on covariance cleaning and nonlinear shrinkage has delivered asymptotically optimal analytical solutions for large covariance matrices. The same framework has been generalized to empirical cross–covariance matrices, whose singular value decomposition identifies canonical comovement modes between two asset sets, with singular values quantifying the strength of each mode and providing natural targets for shrinkage. Existing analytical cross-covariance cleaners are derived under strong stationarity and large-sample assumptions, and they typically rely on mesoscopic regularity conditions such as bounded spectra; macroscopic common modes (e.g., a global market factor) violate these conditions. When applied to real equity returns, where dependence structures drift over time and global modes are prominent, we find that these theoretically optimal formulas do not translate into robust out-of-sample performance. We address this gap by designing a random-matrix-inspired neural architecture that operates in the empirical singular-vector basis and learns a nonlinear mapping from empirical singular values to their corresponding cleaned values. By construction, the network can recover the analytical solution as a special case, yet it remains flexible enough to adapt to non-stationary dynamics and mode-driven distortions. Trained on a long history of equity returns, the proposed method achieves a more favorable bias–variance trade-off than purely analytical cleaners and delivers systematically lower out-of-sample cross–covariance prediction errors. Our results demonstrate that combining random-matrix theory with machine learning makes asymptotic theories practically effective in realistic time-varying markets.
@@ 1 Introduction
Related denoising results exist for additive-noise models [17, 11], but sample (cross-)covariances arise from products of observations and therefore call for a multiplicative-noise cleaning framework. To the best of our knowledge, the only available analytical characterization of an asymptotically optimal singular-value shrinkage for empirical cross-covariance matrices under multiplicative noise is due to Benaych-Georges, Bouchaud, and Potters (BBP) [2]. In the high-dimensional asymptotic regime, they derive a Rotationally Invariant Estimator (RIE) that preserves the empirical singular vectors of $\widehat{\mathbf{\Sigma}}_{XY}$ while shrinking its singular values so as to minimize the asymptotic Mean Squared Error (MSE) relative to the true cross-covariance $\mathbf{\Sigma}_{XY}$. Their result is presented in two versions: a baseline setting with identity marginals ($\mathbf{\Sigma}_{XX}=\mathbf{I}_{n_{x}}$ and $\mathbf{\Sigma}_{YY}=\mathbf{I}_{n_{y}}$), which also covers the practically relevant case of whitened variables, and a more general setting that incorporates non-trivial marginal covariances into the optimal correction. The theory is validated through simulations and controlled benchmarks, but their work does not consider real financial data and relies on strong stationarity assumptions that are rarely met in practice.
@@
In this work, we introduce a physics-informed learning framework for cross–covariance estimation that preserves the rotational invariances of RMT while operating beyond the stationarity and boundedness regimes required by existing analytical shrinkage. The key idea is to parameterize the cleaned estimator directly in the empirical singular-vector basis and to learn a dimension-agnostic nonlinear map from empirical singular values and two-stream marginal projections to cleaned singular values, so that the BBP correction is recovered as a limiting case, but additional degrees of freedom can accommodate macroscopic modes and time variation. Empirically, we observe that the analytically optimal BBP shrinkage can lose robustness on equity returns as the universe size increases in the presence of a dominant market mode, whereas the proposed symmetry-preserving estimator remains stable and yields lower OOS reconstruction error.
@@ 3 Validation on Synthetic and Market Data
Panel A: Finite-Rank Spiked Models – Varying Spike % $\boldsymbol{\xi}$
$\xi=40\%$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(2.46\pm 0.01)\times 10^{-4}}$ $\mathbf{(2.46\pm 0.01)\times 10^{-4}}$ $(2.53\pm 0.01)\times 10^{-4}$
$\xi=30\%$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(1.92\pm 0.01)\times 10^{-4}}$ $\mathbf{(1.92\pm 0.01)\times 10^{-4}}$ $(1.94\pm 0.01)\times 10^{-4}$
$\xi=20\%$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(1.34\pm 0.01)\times 10^{-4}}$ $\mathbf{(1.34\pm 0.01)\times 10^{-4}}$ $\mathbf{(1.34\pm 0.01)\times 10^{-4}}$
$\xi=10\%$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(7.00\pm 0.04)\times 10^{-5}}$ $\mathbf{(7.01\pm 0.04)\times 10^{-5}}$ $(7.10\pm 0.04)\times 10^{-5}$
$\xi=0\%$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(7.87\pm 0.63)\times 10^{-8}}$ $(2.67\pm 0.12)\times 10^{-7}$ $(5.72\pm 0.01)\times 10^{-6}$
@@
Panel B: Heavy-Tailed Bulk Models – Varying Tail Exponent $\boldsymbol{\alpha}$
Gaussian $(2.00\pm 0.00)\times 10^{-3}$ $(6.16\pm 0.00)\times 10^{-4}$ $(6.18\pm 0.00)\times 10^{-4}$ $\mathbf{(6.16\pm 0.00)\times 10^{-4}}$
$\alpha=5.0$ $(2.00\pm 0.00)\times 10^{-3}$ $(6.17\pm 0.00)\times 10^{-4}$ $(6.18\pm 0.00)\times 10^{-4}$ $\mathbf{(6.16\pm 0.00)\times 10^{-4}}$
$\alpha=2.5$ $(2.00\pm 0.00)\times 10^{-3}$ $(6.16\pm 0.00)\times 10^{-4}$ $(6.18\pm 0.00)\times 10^{-4}$ $\mathbf{(6.16\pm 0.00)\times 10^{-4}}$
$\alpha=1.5$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(6.19\pm 0.01)\times 10^{-4}}$ $(6.21\pm 0.01)\times 10^{-4}$ $\mathbf{(6.19\pm 0.01)\times 10^{-4}}$
@@
Panel C: White Heavy-Tailed Models – Varying Tail Exponent $\boldsymbol{\alpha}$
Gaussian $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(5.64\pm 0.01)\times 10^{-4}}$ $(5.65\pm 0.01)\times 10^{-4}$ $\mathbf{(5.64\pm 0.01)\times 10^{-4}}$
$\alpha=5.0$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(5.65\pm 0.01)\times 10^{-4}}$ $(5.66\pm 0.01)\times 10^{-4}$ $\mathbf{(5.65\pm 0.01)\times 10^{-4}}$
$\alpha=2.5$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(2.58\pm 0.08)\times 10^{-4}}$ $\mathbf{(2.59\pm 0.08)\times 10^{-4}}$ $\mathbf{(2.59\pm 0.08)\times 10^{-4}}$
$\alpha=1.5$ $(2.00\pm 0.00)\times 10^{-3}$ $\mathbf{(4.15\pm 0.16)\times 10^{-5}}$ $\mathbf{(4.13\pm 0.16)\times 10^{-5}}$ $\mathbf{(4.37\pm 0.16)\times 10^{-5}}$
@@
Panel D: Gaussian Bulk Models with Common Mode – Varying Mode $\boldsymbol{m}$
$m=0.5$ $(1.13\pm 0.02)\times 10^{-3}$ $(7.93\pm 0.04)\times 10^{-2}$ $\mathbf{(7.92\pm 0.23)\times 10^{-4}}$ $\mathbf{(7.89\pm 0.23)\times 10^{-4}}$
$m=0.3$ $(1.67\pm 0.02)\times 10^{-3}$ $(9.95\pm 0.09)\times 10^{-3}$ $\mathbf{(9.94\pm 0.16)\times 10^{-4}}$ $\mathbf{(9.92\pm 0.16)\times 10^{-4}}$
$m=0.1$ $(1.96\pm 0.00)\times 10^{-3}$ $(8.63\pm 0.05)\times 10^{-4}$ $\mathbf{(8.44\pm 0.03)\times 10^{-4}}$ $(8.54\pm 0.03)\times 10^{-4}$
$m=0.05$ $(1.99\pm 0.00)\times 10^{-3}$ $\mathbf{(7.44\pm 0.02)\times 10^{-4}}$ $(7.46\pm 0.02)\times 10^{-4}$ $(7.62\pm 0.01)\times 10^{-4}$
@@
Table 1: Summary of the average MSE for different estimators across three synthetic regimes over 1000 simulations. Bold marks the lowest MSE; ties are bold when a 95% percentile $10,000$ copies bootstrap cannot distinguish them. The errors indicate the half-width of the confidence intervals (see Appendix 7 for the model definitions). Panel A varies the spike percentage $\xi$ (Finite-Rank). Panel B and C vary the tail exponent $\alpha$ (Heavy-Tailed Bulk and White models), where a lower $\alpha$ implies heavier tails. Panel D is a variation of Panel B (Gaussian) with a one-factor mode injected. All simulations are performed with $n_{x}=200$, $n_{y}=350$, $\Delta t_{\textrm{in}}=500$ as in Ref. [2].
@@
Condition Empirical (MLE) Analytical (BBP) Numerical (CV) Neural Network (NN) Oracle
Panel A: Chronological Split – Varying Universe Size $\boldsymbol{n}$
$n=250$ 0.0342(±0.0005)0.0342\ ($\pm$ 0.0005) $\mathbf{0.0303\ (\pm 0.0004)}$ 0.0330(±0.0004)0.0330\ ($\pm$ 0.0004) $\mathbf{0.0295\ (\pm 0.0003)}$ 0.0167(±0.0001)0.0167\ ($\pm$ 0.0001)
$n=500$ 0.0340(±0.0005)0.0340\ ($\pm$ 0.0005) 0.0317(±0.0005)0.0317\ ($\pm$ 0.0005) 0.0328(±0.0003)0.0328\ ($\pm$ 0.0003) $\mathbf{0.0277\ (\pm 0.0004)}$ 0.0166(±0.0002)0.0166\ ($\pm$ 0.0002)
$n=750$ 0.0340(±0.0003)0.0340\ ($\pm$ 0.0003) 0.0473(±0.0007)0.0473\ ($\pm$ 0.0007) 0.0328(±0.0003)0.0328\ ($\pm$ 0.0003) $\mathbf{0.0298\ (\pm 0.0004)}$ 0.0167(±0.0002)0.0167\ ($\pm$ 0.0002)
$n=1000$ 0.0340(±0.0003)0.0340\ ($\pm$ 0.0003) 0.0656(±0.0008)0.0656\ ($\pm$ 0.0008) 0.0328(±0.0004)0.0328\ ($\pm$ 0.0004) $\mathbf{0.0296\ (\pm 0.0004)}$ 0.0167(±0.0001)0.0167\ ($\pm$ 0.0001)
Panel B: Time-Permutation Control – Varying Universe Size $\boldsymbol{n}$
$n=250$ 0.0111(±0.0001)0.0111\ ($\pm$ 0.0001) 0.0111(±0.0001)0.0111\ ($\pm$ 0.0001) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0106\ (\pm 0.0001)}$ 0.0082(±0.0000)0.0082\ ($\pm$ 0.0000)
$n=500$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0224(±0.0002)0.0224\ ($\pm$ 0.0002) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
$n=750$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0456(±0.0004)0.0456\ ($\pm$ 0.0004) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
$n=1000$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0684(±0.0006)0.0684\ ($\pm$ 0.0006) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
Panel C: Chronological Split – Varying Relative Dimension $\boldsymbol{\nu}$
$\nu=0.1$ 0.0340(±0.0003)0.0340\ ($\pm$ 0.0003) 0.0431(±0.0007)0.0431\ ($\pm$ 0.0007) 0.0328(±0.0004)0.0328\ ($\pm$ 0.0004) $\mathbf{0.0295\ (\pm 0.0005)}$ 0.0167(±0.0001)0.0167\ ($\pm$ 0.0001)
$\nu=0.2$ 0.0341(±0.0004)0.0341\ ($\pm$ 0.0004) 0.0601(±0.0008)0.0601\ ($\pm$ 0.0008) 0.0329(±0.0004)0.0329\ ($\pm$ 0.0004) $\mathbf{0.0296\ (\pm 0.0004)}$ 0.0167(±0.0001)0.0167\ ($\pm$ 0.0001)
$\nu=0.3$ 0.0340(±0.0004)0.0340\ ($\pm$ 0.0004) 0.0687(±0.0007)0.0687\ ($\pm$ 0.0007) 0.0328(±0.0003)0.0328\ ($\pm$ 0.0003) $\mathbf{0.0296\ (\pm 0.0004)}$ 0.0166(±0.0002)0.0166\ ($\pm$ 0.0002)
$\nu=0.4$ 0.0341(±0.0003)0.0341\ ($\pm$ 0.0003) 0.0740(±0.0009)0.0740\ ($\pm$ 0.0009) 0.0328(±0.0003)0.0328\ ($\pm$ 0.0003) $\mathbf{0.0297\ (\pm 0.0004)}$ 0.0167(±0.0002)0.0167\ ($\pm$ 0.0002)
Panel D: Time-Permutation Control – Varying Relative Dimension $\boldsymbol{\nu}$
$\nu=0.1$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0401(±0.0004)0.0401\ ($\pm$ 0.0004) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
$\nu=0.2$ 0.0111(±0.0001)0.0111\ ($\pm$ 0.0001) 0.0622(±0.0006)0.0622\ ($\pm$ 0.0006) $\mathbf{0.0105\ (\pm 0.0001)}$ $\mathbf{0.0106\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
$\nu=0.3$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0730(±0.0008)0.0730\ ($\pm$ 0.0008) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
$\nu=0.4$ 0.0112(±0.0001)0.0112\ ($\pm$ 0.0001) 0.0783(±0.0007)0.0783\ ($\pm$ 0.0007) $\mathbf{0.0106\ (\pm 0.0001)}$ $\mathbf{0.0107\ (\pm 0.0001)}$ 0.0083(±0.0000)0.0083\ ($\pm$ 0.0000)
@@
Table 2: OOS MSE for cross-correlation estimators on financial data covering 2017–2024. For each test window, the estimator is trained on an expanding sample from 1995 up to the year immediately preceding the test period, and the reported value averages 1,000 independent runs for each OOS year. The panels report also the Oracle, which is the lower achievable MSE by a RIE (see Eq. (35)). Panels A and C preserve the chronological order between in-sample and OOS windows, whereas panels B and D shuffle dates. Panels A and B vary the total number of assets for $\nu=0.25$, panels C and D vary the relative dimension $\nu$ for $n=1000$. The errors indicate the half-width of a $95\%$ percentile $100,000$ copies bootstrap confidence interval, bold numbers highlight the methods statistically closer to the Oracle.
@@
Table 2 (panel A) reports the OOS MSE as a function of the total number of assets $n$ for a fixed in-sample window ($\Delta t_{\textrm{in}}=500$). The Maximum-Likelihood Estimator (MLE) provides the unfiltered baseline. Across both the $n$- and $\nu$-sweeps, the proposed NN estimator achieves the lowest error. For small universes (e.g., $n=250$), its performance is not statistically distinguishable from BBP. However, as $n$ increases, BBP deteriorates sharply and can even underperform MLE, indicating that the analytical shrinkage can become harmful in this regime. We attribute this failure to the presence of a macroscopic market mode in real returns, which violates the boundedness conditions underlying BBP. This interpretation is supported by two additional diagnostics. First, the numerical CV estimator is not affected by the same degradation and remains comparable to the NN. Second, in Appendix 8.3 we show that pre-processing the data by removing the mode, operationally defined as the daily cross-sectional mean return, eliminates the gross of the BBP underperformance. Finally, when we shuffle the time series to break potential non-stationarity between in-sample and OOS periods, BBP remains unstable, whereas CV continues to track the NN, further indicating that the dominant driver is the mode rather than temporal non-stationarity (see panel B of Table 2). Notably, the NN performance remains robust even for $n>500$, which constitutes a fully out-of-distribution regime for the model. This region lies beyond the matrix sizes observed during training ($n\leq 500$), and it also corresponds to a different historical period than the data used to fit the network. To further probe this extrapolative regime, in panel C of Table 2 we unpack the $n=1000$ case by varying the relative dimension $\nu$, and we find that the NN gain remains stable across the full range of $\nu$, indicating that its advantage is not concentrated in a specific configuration.
@@
Our working hypothesis is that this behavior is precisely the one targeted by the proposed architecture. In stationary settings, the network should primarily act as a noise-cleaning operator on the empirical spectrum. In contrast, when applied to non-stationary financial data, it can interpolate between denoising and a learned predictive correction that exploits persistent structure beyond sampling noise. To confirm that this gain is derived from capturing genuine temporal signals (non-stationarity), we replicate the analysis on temporally shuffled data. Shuffling destroys the serial correlations responsible for the predictive edge, creating a purely stationary environment. Panels C and C of Table 2 illustrate that performance degradation of BBP persists even in the shuffled regime, confirming the underlying structural driver. As expected, the CV estimator does not degrade and instead matches the NN performance indicating that in a stationary regime the NN learn how to denoise only sample noise. This interpretation is further supported by the feasibility diagnostic reported in Appendix 8.4.